if the interior angle of a regular polygon is 2880 degrees find the value of side of one of it exterior
Answer by fractalier(6550)

The sum of the interior angles is 180(n-2)...here that total is 2880, so we have
180(n-2) = 2880
n-2 = 16
n = 18 sides
The external angle is found by 360/n = 360/18 = 20 degrees